# Terahertz-induced second-harmonic generation in quantum paraelectrics: hot-phonon effect

### v1

@@ Abstract
Recent terahertz-pump second-harmonic-generation(SHG)-probe measurements of quantum paraelectrics observed a significant long-lived non-oscillatory SHG component following an ultrafast resonant excitation of the soft mode, which was interpreted as a signature of terahertz-induced transient ferroelectric order. Here we propose a temperature-dependent dynamic model incorporating the hot-phonon effect to simulate the soft-mode behaviors under ultrafast terahertz excitation. Its application to paraelectric KTaO3 produces quantitatively most of the features exhibited in our time-resolved SHG measurements and those in existing literature, including a long-lived non-oscillatory SHG response, SHG oscillations at twice the soft-mode frequency, SHG dampings as well as temperature and field-strength dependencies. We conclude that the observed terahertz-induced non-oscillatory SHG response in quantum paraelectrics is a consequence of the induced nonequilibrium hot-phonon effect, offering an alternative to its existing interpretation as a signature of transient ferroelectric order.
@@
Introduction.—The quantum criticality in condensed matter physics describes the ordering of a quantum phase that occurs at/near zero temperature. This phenomenon has attracted considerable attention, partly due to its distinct characteristics and unexpected physics arising from the low-lying collective excitations. Extensive research over the past few decades has suggested the presence of the quantum criticality in certain strongly-correlated materials, but which are often complex with various intertwined quantum orders. A notable exception is the displacive quantum paraelectrics Müller and Burkard (1979); Rowley et al. (2014); Cowley (1996); Fujishita et al. (2016); Verdi et al. (2023); Yang and Chen (2024), where there is a strong competition between quantum fluctuations and ferroelectric ordering. This class of material is supposed to transition from the paraelectric to ferroelectric states at low temperature due to the lattice dynamical instability Verdi et al. (2023); Wu et al. (2022); Müller and Burkard (1979); Singh (1996); Cowley (1996), but the zero-point lattice vibrations prohibits the long-range ferroelectric order Müller and Burkard (1979); Verdi et al. (2023); Yang and Chen (2024); Wu et al. (2022), leading to an incipient ferroelectricity, sometimes referred to as hidden ferroelectric phase Li et al. (2019); Cheng et al. (2023).
@@
While doping Andrews (1985); Toulouse et al. (1992); Aktas et al. (2014); Rischau et al. (2017) or isotope substitution Takesada et al. (2006); Rowley et al. (2014) can turn quantum paraelectrics to ferroelectrics, transiently reaching the hidden ferroelectric phase through ultrafast manipulation with intense femtosecond-pulsed laser De La Torre et al. (2021); Fechner et al. (2024); Basini et al. (2024); Nova et al. (2019) is particularly appealing, given its promising potential for applications in memory and computational devices Prosandeev et al. (2021, 2022). One strategy is to coherently drive the so-called soft-phonon mode—associated with the lattice dynamical instability Yamada and Shirane (1969); Fleury and Worlock (1968); Sirenko et al. (2000); Shirane et al. (1967)—into the nonlinear regime, aiming to create transient ferroelectric order. For this purpose, several terahertz(THz)-pumped second-harmonic-generation(SHG)-probe measurements have been performed in the protypical displacive quantum paraelectrics, SrTiO3 Li et al. (2019) and KTaO3 single crystals Cheng et al. (2023); Li et al. (2023). At low temperatures, a significant long-lived non-oscillatory SHG component, superimposed by a clear SHG oscillation at twice the soft-mode frequency, was observed after a THz pulse Li et al. (2019); Cheng et al. (2023); Li et al. (2023). The origin and precise mechanisms of this non-oscillatory component and soft-mode frequency doubling phenomenon Li et al. (2019); Cheng et al. (2023); Li et al. (2023) remain unclear Zhuang et al. (2023); Shin et al. (2022). The observed non-oscillatory SHG background is commonly interpreted as a signature of a THz-induced transient ferroelectric order, possibly arising from a THz-driven intrinsic lattice displacement Li et al. (2019), or from a THz-induced long-range correlation between extrinsic local polar structures by defects Cheng et al. (2023). However, such an interpretation seems to contradict several other findings. For example, it was realized that the THz pulses up to 500 kV/cm were insufficient to produce a global intrinsic ferroelectricity Cheng et al. (2023), whereas inducing a long-range correlation between extrinsic defect dipoles via an ultrafast manipulation does not require coherently driving an intrinsic soft mode. More importantly, soft modes are experimentally observed to go through hardening with an increasing THz-field strength Li et al. (2019); Cheng et al. (2023); Li et al. (2023), suggesting an intense THz field drives the quantum paraelectrics away from rather than towards ferroelectricity, since the soft-mode hardening in displacive paraelectrics is an indicator of departure from ferroelectricity.
@@
On the other hand, similar non-oscillatory components in pump-probe measurements have been commonly observed in other subfields, e.g., in measurements of the interband transitions in semiconductors (e.g., graphene Hafez et al. (2018)) and collective excitations in superconductors Matsunaga et al. (2013, 2014); Shimano and Tsuji (2020); Yang and Wu (2018); Cui et al. (2019), and they were attributed to the induced nonequilibrium hot-quasiparticle effect. For example, in superconductors, by using intense THz pulses, one can resonantly excite the amplitude mode of the superconducting order parameter Matsunaga et al. (2013, 2014); Shimano and Tsuji (2020); Yang and Wu (2019, 2023), namely the Higgs-mode excitation Pekker and Varma (2015). In most of the measurements, a non-oscillatory component was observed and can persist for a long time after the THz pulse Matsunaga et al. (2013, 2014); Shimano and Tsuji (2020), as a consequence of the nonequilibrium hot-quasiparticle effect Matsunaga et al. (2013, 2014); Shimano and Tsuji (2020); Yang and Wu (2018); Cui et al. (2019). In contrast, such a hot-quasiparticle effect in ultrafast-optical responses of ferroelectrics and paraelectrics has largely been overlooked.
In this work, we propose a temperature-dependent dynamic model to describe the soft-mode behaviors under an ultrafast resonant excitation by incorporating the hot-phonon effect, and examine the THz-induced SHG responses in displacive quantum paraelectric KTaO3 using a combination of theoretical and experimental studies. Numerical simulations based on the developed dynamic model quantitatively produce all the main features exhibited in our time-resolved SHG measurements, including a long-lived non-oscillatory SHG response, SHG oscillations at twice the soft-mode frequency, SHG dampings as well as temperature and field-strength dependence. We therefore attribute the THz-induced non-oscillatory SHG component to the nonequilibrium hot-phonon effect, offering an alternative to its existing interpretation as a signature of transient ferroelectric order in quantum paraelectrics Li et al. (2019); Cheng et al. (2023); Li et al. (2023).
@@
Furthermore, we also explore the SHG responses of a ferroelectric KTaO3 after an ultrafast THz excitation to understand the individual responses of actual ferroelectric nano-regions (e.g., extrinsic local polar structures by defects in quantum paraelectrics Cheng et al. (2023)). In this case, both our theoretical simulations and experimental measurements show a THz-induced long-lived SHG oscillation at the single polar-mode frequency without any evident signature for the non-oscillatory component, in contrast to the observed SHG response of quantum paraelectrics under an ultrafast THz excitation, suggesting that the previously reported resonant SHG features in quantum paraelectrics does not come from ferroelectric nano-regions.
@@
Experimental setup.—We use intense single-cycle THz pump pulses (up to 210210210~{}210kV/cm), generated from the optical rectification in a LiNbO3 prism Li et al. (2023), to resonantly excite the soft mode in a quantum paraelectric KTaO3 single crystal and detect the time-resolved SHG signal via a femtosecond optical-frequency (800-nm) probe pulse as a function of delay time with respect to the pump pulse. The field directions of the THz pump and the optical-frequency probe pulses are both set along the [100] direction in KTaO3 crystal. The time-resolved signal is measured using a GaP photodiode and a blue filter after the KTaO3 sample, which eliminates the fundamental 800-nm wavelength from the detected light. The waveform of the employed single-cycle THz pump pulses is shown in Fig. 1(b). More details can be found in Supplemental Materials.
@@
where $b$ is an anharmonic coefficient, related to three-soft-phonon interactions; $m_{p}$ denotes the effective mass Sivasubramanian et al. (2004); ${\bf E}(t)$ represents the THz field; $\alpha(T)$ is the harmonic coefficient, and using the self-consistent renormalization theory within the path-integral approach Yang and Chen (2024), the equilibrium harmonic coefficient is derived as $\alpha_{e}(T)=\alpha_{e}(T=0)+{b}{C}(T)$ (see Supplementary Materials). Here, $C(T)$ requires a self-consistent formulation of the bosonic thermal excitation of the soft phonons:
@@
Here, we have introduced a damping term, with $\gamma$ being the damping rate. This damping should be dominated by three-phonon scattering between two soft phonons and one acoustic phonon, leading to a $T$-dependent $\gamma(T)$. In principle, the evolution of $\alpha(t)$ should incorporate a fully microscopic bosonic Boltzmann equation of the soft phonons. However, such a microscopic treatment is complex and will not change the main results and conclusions in this work. Thus, here we employ the Allen-Cahn-like relaxation equation extensively used in the phase-field method Chen (2002) (see Supplemental Materials):
@@
where we take $\eta<0$ since the soft-phonon number increases after the excitation and leads to the increase of $\alpha(t)$ according to Eq. (2); $\tau_{E}$ is the energy-relaxation time of the system.
For quantum paraelectrics, we first self-consistently solve the coupled Eqs. (2) and (3) using only the knowledge about the ground-state parameters $\alpha_{e}(T=0)$ and $b$, to obtain the equilibrium $\alpha_{e}(T)>0$, and hence, $P_{0}^{2}\equiv 0$. We then solve the dynamic equations [Eqs. (4) and (5)] using the experimental waveform of the single-cycle THz pump pulse [Fig. 1(b)] as the input field ${\bf E}(t)$, resulting in the temperature-dependent dynamics of the soft mode under an ultrafast excitation.
It has been established that the inverse dielectric function $1/\varepsilon(T)\propto\alpha(T)$ in quantum paraelectrics Rowley et al. (2014); Yang and Chen (2024), thereby leading to the THz-induced SHG intensity $\delta{\rm I}_{\rm SHG}(t)\propto\delta\alpha(t)$ (See Supplemental Materials). In specific simulations, we consider the damping rate $\gamma(T)$ as the only fitting parameter for temperature variation in the experimental measurements. Other parameters used in simulations are based on several independent experimental measurements (See Supplementary Materials).
@@
Figure 1: (a) Schematic illustration of the generation (heating process) and decay (cooling process) of the hot-phonon effect during a nonlinear ultrafast resonant excitation, which creates a long-lived non-oscillatory SHG component. (b) The waveform and (c) waveform squared of the single-cycle THz pump pulse employed in experiments and simulations. (d) Theoretically calculated and (e) experimentally measured time-resolved SHG in paraelectric KTaO3 at different $T$.
@@
Figure 2: (a) Detailed comparison between theoretical and experimental results of the time-resolved SHG signal. (b) Fourier transform of the measured SHG oscillations at different temperatures. It is noted that at temperatures below 100 K, besides the widely reported soft mode that emerges around 1.8 THz, there exists another mode around lower frequency of 0.4 THz, which should correspond to a collective vector mode emerged in displacive ferroelectrics and quantum paraelectrics as proposed in Ref. Yang and Chen (2024). (c) Theoretical and experimental results of the soft-mode excitation gap. The inset shows the employed damping rate $\gamma(T)$ (squares) in our simulation, and a fitting curve: $\gamma(T)\propto 1+2/[\exp({T_{\rm ac}/T})-1]$ with $T_{\rm ac}=76~{}$K. (d) SHG peaks as a function of THz-field power at 20 K. The inset shows the numerical time-resolved SHG signal at different field strengths.
For an analytical analysis of the THz excitation, we assume a pump field in the single-frequency form: ${\bf E}(t)\approx{{\bf E}_{0}}\cos{(\Omega{t})}$, with $\Omega$ being the THz-field frequency. From Eqs. (4) and (5), neglecting all damping terms, one approximately has $\partial_{t}\delta\alpha(t)\approx\frac{|\eta|{E}^{2}_{0}\cos^{2}(\Omega{t})}{2[m_{p}(\omega^{2}_{q=0}-\Omega^{2})]^{2}}$ (see Supplementary Materials), leading to the THz-induced dynamic behavior of $\delta\alpha(t)$:
@@
Our numerical and experimental results for the SHG signal peaks as a function of the THz-field power are plotted in Fig. 2(d). Clearly, they are proportional to the THz-field power, i.e., $\delta{\rm I}_{\rm SHG}\propto{E^{2}_{\rm pump}E^{4}_{\rm probe}}$, as also observed in the previous measurements Li et al. (2019); Cheng et al. (2023); Li et al. (2023) and in agreement with the analysis above in Eq. (6). This suggests that the THz-induced SHG response in quantum paraelectrics is a second-order response to the THz pump field, a near-equilibrium response, rather than the intricate far-from-equilibrium dynamics that usually manifests at higher-order responses, and there is no generation of a transient ferroelectric order. Actually, it should be emphasized that a finite SHG signal detected by the 800-nm probe field with frequency of 1.55 eV in fact can not justify the emergence of ferroelectric order (see Supplementary Materials).
@@
Finally, it is noted that to explain the observed long-lived non-oscillatory SHG component, Cheng et al. has proposed a potential origin in Ref. Cheng et al. (2023), a THz-induced correlation between the local polar structures (i.e., ferroelectric nano-regions) that arise from the extrinsic defects, leading to a global ferroelectric-like response. To examine this possibility, we explore the THz-induced SHG response of a ferroelectric KTaO3, experimentally obtained through an annealing process under high vacuum FEK . We then conducted the THz-pump SHG-probe measurements on the ferroelectric KTaO3 crystal, and the results are plotted in Fig. 3. As shown in Fig. 3, the ferroelectric KTaO3 exhibits a totally different THz-induced SHG responses from the quantum paraelectric KTaO3. The THz-field strength dependence of the SHG response (inset of Fig. 3) shows a clear butterfly-shape hysteresis loop, which demonstrates the presence of the ferroelectricity. As for the time-resolved SHG signal (Fig. 3), a clear oscillation develops after the THz pulse ($t>2~{}$ps) and persists up to 7 ps, suggesting a coherent/resonant driving of a collective excitation (i.e., a polar mode). However, no evident non-oscillatory component was observed, and no polar-mode frequency doubling phenomenon occurs in the resonant excitation as the observed SHG response during the pump pulse ($t<2$ ps) just follows the pump-pulse waveform [Fig. 1(b)]. This suggests that the THz-induced SHG signal in ferroelectric KTaO3 is a linear response to the THz pump field, i.e., $\delta{\rm I}_{\rm SHG}\propto{E_{\rm pump}}{E^{4}_{\rm probe}}$ as a consequence of the breaking of the global lattice inversion symmetry by the existing ferroelectric order. These resonant features in the ferroelectric state are in contrast to the ones in the quantum paraelectric state, and therefore, it should suggest that the ferroelectric nano-regions are not the origin for the observed SHG resonant features in the quantum paraelectrics.
@@
with $\Delta=\sqrt{-2\alpha_{e}/m_{p}}$ being the polar-mode excitation gap. Consequently, the presence of the ferroelectric order leads to the emergence of a linear response (first term) to the THz field, in agreement with our numerical simulations and experimental measurements as well as symmetry analysis. This response dominates the THz-induced SHG signal at a relatively weak THz field, whereas the hot-phonon effect (non-oscillatory component) and soft-mode frequency doubling phenomenon that manifest in second-order excitation, i.e., second term, are masked in the THz-induced SHG response.
In summary, by combining numerical simulations based on a developed dynamic model and experimental measurements on the THz-induced time-resolved SHG responses in paraelectric KTaO3, we conclude that the observed long-lived non-oscillatory component under an ultrafast excitation in recent THz-pump SHG-probe experiments Li et al. (2019); Cheng et al. (2023) is a consequence of the induced nonequilibrium hot-phonon effect. Actually, based on this understanding of the hot-phonon effect, the previously observed soft-mode hardening as THz-field strength increases Li et al. (2019); Cheng et al. (2023); Li et al. (2023) can also be understood: the increase in the field strength enhances the hot-phonon effect and promotes the soft-phonon temperature during the nonequilibrium process, thereby leading to a soft-mode hardening.
Acknowledgements.
F.Y. and X.J.L. did the theoretical and experimental studies, respectively, and contributed equally to this work. F.Y. and L.Q.C. acknowledge support from the US Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145 as part of Computational Materials Sciences Program. F.Y. and L.Q.C. also appreciate the generous support from the Donald W. Hamer Foundation through a Hamer Professorship at Penn State. X.J.L. and D.T. acknowledge support by the National Science Foundation under Grant No. DMR-1554866.

### v2

@@ Abstract
Recent THz-pump second-harmonic-generation(SHG)-probe measurements of quantum paraelectrics observed a significant long-lived non-oscillatory SHG component following an ultrafast resonant excitation of the soft mode, which was interpreted as a signature of THz-induced transient ferroelectric order. We propose that the THz-induced modulation of the SHG signal can be attributed solely to the dynamic variation of the dielectric environment associated with the lattice background, which reflects the coherent response of soft mode under THz pumping. We develop a temperature-dependent dynamic model incorporating the hot-phonon effect to simulate the soft-mode behaviors under ultrafast THz excitation. Its application to paraelectric KTaO3 produces quantitatively most of the features exhibited in our time-resolved SHG measurements and those in existing literature, including a long-lived non-oscillatory SHG response, SHG oscillations at twice the soft-mode frequency, SHG dampings as well as temperature and field-strength dependencies. We conclude that the observed THz-induced non-oscillatory SHG response in quantum paraelectrics is a consequence of the nonequilibrium hot-phonon effect, offering an alternative to its existing interpretation as a signature of transient ferroelectric order.
@@
Introduction.—The quantum criticality in condensed matter physics describes the ordering of a quantum phase that occurs near zero temperature. This phenomenon has attracted considerable attention, due to its distinct characteristics arising from the low-lying collective excitations. Extensive research over past few decades has suggested the presence of quantum criticality in strongly-correlated materials, which are often complex with various intertwined quantum orders. A notable exception is the displacive quantum paraelectrics Müller and Burkard (1979); Rowley et al. (2014); Cowley (1996); Fujishita et al. (2016); Verdi et al. (2023); Yang and Chen (2024); Singh (1996); Wu et al. (2022), where a strong competition between quantum fluctuations and ferroelectric ordering exists. This class of material is supposed to transition from the paraelectric to ferroelectric states at low temperature due to lattice dynamical instability, but the zero-point lattice vibrations prohibits the long-range ferroelectric order Müller and Burkard (1979); Verdi et al. (2023); Yang and Chen (2024); Wu et al. (2022), leading to an incipient ferroelectricity, sometimes referred to as hidden ferroelectric phase Li et al. (2019); Cheng et al. (2023).
@@
While doping Andrews (1985); Toulouse et al. (1992); Aktas et al. (2014); Rischau et al. (2017) or isotope substitution Takesada et al. (2006); Rowley et al. (2014) can turn quantum paraelectrics to ferroelectrics, transiently reaching the hidden ferroelectric phase through ultrafast manipulation with intense femtosecond-pulsed laser De La Torre et al. (2021); Fechner et al. (2024); Basini et al. (2024); Nova et al. (2019) is particularly appealing Prosandeev et al. (2021, 2022). One strategy is to coherently drive the so-called soft-phonon mode—associated with lattice dynamical instability Yamada and Shirane (1969); Fleury and Worlock (1968); Sirenko et al. (2000); Shirane et al. (1967)—into the nonlinear regime, aiming to create transient ferroelectric order. Several terahertz(THz)-pumped second-harmonic-generation(SHG)-probe measurements have therefore been performed in typical displacive quantum paraelectrics, SrTiO3 Li et al. (2019) and KTaO3 single crystals Cheng et al. (2023); Li et al. (2023). At low temperatures, a significant long-lived non-oscillatory SHG component, superimposed by a clear SHG oscillation at twice the soft-mode frequency, was observed after a THz pulse Li et al. (2019); Cheng et al. (2023); Li et al. (2023). The origin and precise mechanisms of this non-oscillatory component and soft-mode frequency doubling phenomenon Li et al. (2019); Cheng et al. (2023); Li et al. (2023) remain unclear Zhuang et al. (2023); Shin et al. (2022). The observed non-oscillatory SHG background is commonly interpreted as a signature of a THz-induced transient ferroelectric order, possibly arising from a THz-driven intrinsic lattice displacement Li et al. (2019), or from a THz-induced long-range correlation between extrinsic local polar structures by defects Cheng et al. (2023). However, such an interpretation seems to contradict several other findings. For example, it was realized that the THz pulses up to 500 kV/cm were insufficient to produce a global intrinsic ferroelectricity Cheng et al. (2023), whereas inducing a long-range correlation between extrinsic defect dipoles via an ultrafast manipulation does not require coherently driving an intrinsic soft mode. More importantly, soft modes are experimentally observed to go through hardening with an increasing THz-field strength Li et al. (2019); Cheng et al. (2023); Li et al. (2023), suggesting an intense THz field drives the quantum paraelectrics away from rather than towards ferroelectricity, since the soft-mode hardening in displacive paraelectrics is an indicator of departure from ferroelectricity.
@@
On the other hand, similar non-oscillatory components in pump-probe measurements have been commonly observed in other subfields, e.g., in measurements of the interband transitions in semiconductors (e.g., graphene Hafez et al. (2018)) and collective excitations in superconductors Matsunaga et al. (2013, 2014); Shimano and Tsuji (2020); Yang and Wu (2018); Cui et al. (2019), and they were attributed to the induced nonequilibrium hot-quasiparticle effect. For example, in superconductors, by using intense THz pulses, one can resonantly excite the amplitude mode of the superconducting order parameter Matsunaga et al. (2013, 2014); Shimano and Tsuji (2020); Yang and Wu (2019, 2023), namely the Higgs-mode excitation Pekker and Varma (2015). In most of the measurements, a non-oscillatory component was observed and can persist for a long time after the THz pulse Matsunaga et al. (2013, 2014); Shimano and Tsuji (2020), as a consequence of the nonequilibrium hot-quasiparticle effect Matsunaga et al. (2013, 2014); Shimano and Tsuji (2020); Yang and Wu (2018); Cui et al. (2019). The hot-quasiparticle effect is a fundamental aspect of nonequilibrium dynamics, applicable to both fermions (electrons) and bosons (phonons) under an external excitation due to the energy input. However, a dedicated hot-phonon effect for purely soft-phononic dynamics under ultrafast optical excitation in ferroelectrics and paraelectrics has largely been overlooked.
In this work, we propose that a finite probe-field SHG signal originates from the oxygen-vacancy defects and remains nearly unchanged during the THz excitation. The observed THz-induced SHG modulation can be attributed solely to the dynamic variation of the dielectric environment associated with the lattice background, which reflects the soft mode’s coherent response under THz pumping. For this purpose, we develop a temperature-dependent dynamic model to describe the soft-mode behaviors under an ultrafast resonant excitation by incorporating the hot-phonon effect, and examine the THz-induced SHG responses in quantum paraelectric KTaO3 using a combination of theoretical and experimental studies.
@@
Simulations based on our model quantitatively produce all the main features exhibited in our time-resolved SHG measurements, including a long-lived non-oscillatory response, oscillations at twice the soft-mode frequency, SHG dampings as well as temperature and field-strength dependence. We therefore attribute the THz-induced non-oscillatory SHG component in quantum paraelectrics to the nonequilibrium hot-phonon effect. Additionally, we also explore the SHG responses of a ferroelectric KTaO3 after an ultrafast THz excitation to understand the individual responses of actual ferroelectric nano-regions (e.g., extrinsic local polar structures by defects Cheng et al. (2023)). In this case, both our theoretical simulations and experimental measurements show a THz-induced long-lived SHG oscillation at the single polar-mode frequency without any evident signature for the non-oscillatory component, in contrast to the observed THz-induced SHG response of quantum paraelectrics, suggesting that the previously reported resonant SHG features in quantum paraelectrics does not come from ferroelectric nano-regions.
@@
Experimental setup.—We use intense single-cycle THz pump pulses (up to 210210210~{}210kV/cm), generated from the optical rectification in a LiNbO3 prism Li et al. (2023), to resonantly excite the soft mode in a quantum paraelectric KTaO3 single crystal and detect the time-resolved SHG signal via a femtosecond optical-frequency (800-nm) probe pulse as a function of delay time with respect to the pump pulse. The field directions of the THz pump and the optical-frequency probe pulses are both set along the [100] direction in KTaO3 crystal. The time-resolved SHG signal is measured using a blue filter after the sample, which eliminates the fundamental 800-nm wavelength from the detected light. The waveform of the employed single-cycle THz pump pulses is shown in Fig. 1(b). For our experimental conditions (for example, at mediate 80 K, and a laser spot size around 50 $\mu$m focused by a 200 mm lens), the SHG signal read on a lock-in amplifier detected by a GaP photodiode detector is approximately 2.9 $\mu$V with 220 $\mu$W probe intensity, confirming negligible static contribution. Under similar experimental condition but 500 $\mu$W probe intensity, after resonant excitation by a 210 kV/cm THz field, the reading of the SHG signal reaches up to approximately 120 $\mu$V at the peak position in the positive time delay, reflecting a significant (8-times) enhancement beyond static defect-related background.
@@
where $b$ is an anharmonic coefficient, related to three-phonon interactions; $m_{p}$ denotes the effective mass Sivasubramanian et al. (2004); ${\bf E}(t)$ represents the THz field; $\alpha(T)$ is the harmonic coefficient, and using the self-consistent renormalization theory within the path-integral approach Yang and Chen (2024), the equilibrium harmonic coefficient is derived as $\alpha_{e}(T)=\alpha_{e}(T=0)+{b}{C}(T)$ (see Supplemental Materials sup ). Here, $C(T)$ requires a self-consistent formulation of the bosonic thermal excitation of the soft phonons:
@@
Here, we have introduced a damping term, with $\gamma$ being the damping rate. This damping should be dominated by three-phonon scattering between two soft phonons and one acoustic phonon, leading to a $T$-dependent $\gamma(T)$. In principle, the evolution of $\alpha(t)$ should incorporate a fully microscopic bosonic Boltzmann equation of the soft phonons. However, such a microscopic treatment is complex and will not change the main results and conclusions in this work. Thus, here we employ the Allen-Cahn-like relaxation equation extensively used in the phase-field method Chen (2002) (see Supplemental Materials sup ):
@@
where we take $\eta<0$ since the soft-phonon number increases after the excitation and leads to the increase of $\alpha(t)$ according to Eq. (2); $\tau_{E}$ is the energy-relaxation time of the system. The physical picture of hot-phonon effect is shown in Fig. 1(a). The initial THz field induces a hot-phonon effect due to the energy input, which manifests itself as the increase of the harmonic coefficient of the potential well, leading to an increased nonoscillatory SHG component in the first picosecond and hence the hardening of the soft mode oscillation (green arrow). This effect gradually decays after the THz field due to cooling via interactions with acoustic phonons, leading to exponential decay of the non-oscillatory SHG component.
For the simulation in quantum paraelectrics, we first self-consistently solve the coupled Eqs. (2) and (3) using only the knowledge about the ground-state parameters $\alpha_{e}(T=0)$ and $b$, to obtain the equilibrium $\alpha_{e}(T)>0$, and hence, $P_{0}^{2}\equiv 0$. We then solve the dynamic equations [Eqs. (4) and (5)] using experimental waveform of the single-cycle THz pump pulse [Fig. 1(b)] as the input field ${\bf E}(t)$, resulting in the temperature-dependent soft-mode dynamics under an ultrafast excitation.
The inverse dielectric function $1/\varepsilon(T)\propto\alpha(T)$ in quantum paraelectrics Rowley et al. (2014); Yang and Chen (2024), leading to the THz-induced SHG intensity $\delta{\rm I}_{\rm SHG}(t)\propto\delta\alpha(t)$ (See Supplemental Materials sup ). In specific simulations, we consider the damping rate $\gamma(T)$ as the only fitting parameter for temperature variation in the experimental measurements. Other parameters used in simulations are based on several independent experimental measurements.
Figure 1: (a) Schematic illustration of the generation (heating process) and decay (cooling process) of the hot-phonon effect during a nonlinear ultrafast resonant excitation. (b) The waveform and (c) waveform squared of the single-cycle THz pump pulse employed in experiments and simulations. (d) Theoretically calculated and (e) experimentally measured time-resolved SHG in paraelectric KTaO3 at different $T$.
@@
Figure 2: (a) Detailed comparison between theoretical and experimental results in Fig. 1(d) and (e). (b) Fourier transform of the measured SHG oscillations at different $T$. At temperatures below 100 K, besides the widely reported soft mode that emerges around 1.8 THz, there emerges another mode around lower frequency of 0.4 THz, which should correspond to a collective vector mode emerged in displacive ferroelectrics and quantum paraelectrics as proposed in Ref. Yang and Chen (2024). (c) Theoretical and experimental results of the soft-mode excitation gap. The inset shows the employed damping rate $\gamma(T)$ (squares) in our simulation, and a fitting curve mentioned in main text with $T_{\rm ac}=76~{}$K. (d) SHG peaks versus THz-field power at 20 K. The inset shows the numerical time-resolved SHG signal at different field strengths.
@@
For an analytical analysis of the THz excitation, we assume a pump field in the single-frequency form: ${\bf E}(t)\approx{{\bf E}_{0}}\cos{(\Omega{t})}$, with $\Omega$ being the THz-field frequency. From Eqs. (4) and (5), neglecting all damping terms, one approximately has $\partial_{t}\delta\alpha(t)\approx\frac{|\eta|{E}^{2}_{0}\cos^{2}(\Omega{t})}{2[m_{p}(\omega^{2}_{q=0}-\Omega^{2})]^{2}}$ (see Supplemental Materials sup ), leading to the THz-induced dynamic behavior of $\delta\alpha(t)$:
@@
Our numerical and experimental results for the SHG signal peaks as a function of the THz-field power are plotted in Fig. 2(d). They are proportional to the THz-field power, i.e., $\delta{\rm I}_{\rm SHG}\propto{E^{2}_{\rm pump}E^{4}_{\rm probe}}$, as also observed in previous measurements Li et al. (2019); Cheng et al. (2023); Li et al. (2023) and in agreement with the analysis above. This suggests that the THz-induced SHG response in quantum paraelectrics is a second-order response to the THz field.
@@
Finally, it is noted that to explain the observed long-lived non-oscillatory SHG component, Cheng et al. has proposed a potential origin in Ref. Cheng et al. (2023), a THz-induced correlation between the local polar structures (i.e., ferroelectric nano-regions) that arise from the extrinsic defects, leading to a global ferroelectric-like response. To examine this possibility, we explore the THz-induced SHG response of a ferroelectric KTaO3, experimentally obtained through an annealing process FEK . We then conducted the THz-pump SHG-probe measurements on the ferroelectric KTaO3 crystal, and the results are plotted in Fig. 3. As shown in Fig. 3, the ferroelectric KTaO3 exhibits a totally different THz-induced SHG responses from the quantum paraelectric KTaO3. The THz-field strength dependence of the SHG response (inset of Fig. 3) shows a butterfly-shape hysteresis loop of the THz-induced SHG change not , suggesting the presence of the ferroelectricity. As for the time-resolved SHG signal (Fig. 3), a clear oscillation develops after the THz pulse ($t>2~{}$ps) and persists up to 7 ps, suggesting a coherent/resonant driving of a collective excitation (i.e., a polar mode). However, no evident non-oscillatory component was observed, and no polar-mode frequency doubling phenomenon occurs in the resonant excitation as the observed SHG response during the pump pulse ($t<2$ ps) just follows the pump-pulse waveform [Fig. 1(b)]. This suggests that the THz-induced SHG signal in ferroelectric KTaO3 is a linear response to the THz pump field, i.e., $\delta{\rm I}_{\rm SHG}\propto{E_{\rm pump}}{E^{4}_{\rm probe}}$ as a consequence of the breaking of the global lattice inversion symmetry by the existing ferroelectric order. These resonant features in the ferroelectric state are in contrast to the ones in the quantum paraelectric state, and therefore, it should suggest that the ferroelectric nano-regions are not the origin for the observed SHG resonant features in the quantum paraelectrics.
@@
with $\Delta=\sqrt{-2\alpha_{e}/m_{p}}$ being the polar-mode excitation gap. The first term of this expression corresponds to the linear response of the ferroelectric order to the THz field. The presence of the ferroelectric order leads to the emergence of a linear response to the THz field, and it dominates the THz-induced SHG signal at a relatively weak THz field, in agreement with our numerical simulations and symmetry analysis as well as our observations in experimental measurements on annealed, ferroelectric KTaO3 here, where the SHG intensity follows the THz pulse waveform and scales linearly with the pump field. The hot-phonon effect (non-oscillatory component) and soft-mode frequency doubling that manifest in second-order excitation (observed in quantum paraelectrics) are masked by the dominant linear response, which is why these features are not observed in our measurements on ferroelectric KTaO3, leading to the appearance of polar-mode oscillations, a linear dependence of the THz-induced SHG on pump field, and the absence of a non-oscillatory SHG component unless a significant high-order excitation by strongly intense field is present. Actually, similar phenomena have been widely observed in pump-probe measurements of other ferroelectric materials von Hoegen et al. (2018); Grishunin et al. (2017); Miyamoto et al. (2018); Mishina et al. (2018), including the hysteresis loop of the THz-induced SHG change Grishunin et al. (2017); Mishina et al. (2018).
In summary, combining numerical simulations and experimental measurements on the THz-induced time-resolved SHG responses in paraelectric KTaO3, we conclude that the observed long-lived non-oscillatory component in previous experiments Li et al. (2019); Cheng et al. (2023) is a result of nonequilibrium hot-phonon effect. Then, the observed soft-mode hardening as THz-field strength increases Li et al. (2019); Cheng et al. (2023); Li et al. (2023) can also be understood: the increase in the field strength enhances the hot-phonon effect and promotes the soft-phonon temperature during the nonequilibrium process, thereby leading to a soft-mode hardening.
Previous experiments Li et al. (2019); Cheng et al. (2023) assumes that a finite SHG signal necessarily indicates the emergence of ferroelectric order, which breaks inversion symmetry. However, we propose and experimentally demonstrate that while a finite SHG signal indicates the inversion symmetry breaking, it does not necessarily imply the presence of ferroelectric order. Our measurements (refer to Supplemental Material sup ) find that increasing oxygen vacancies in paraelectric KTaO3 can significantly enhance the SHG intensity (even without the THz pump) Der Horst et al. (1996), while no signs of ferroelectric order are present in the system. This indicates that a finite probe-field SHG polarization can exist in quantum paraelectrics due to inevitable oxygen vacancies. It most likely arises from the two-photon inter-band transitions of electrons, mediated by the electronic defect states associated with the oxygen vacancies, which break translational symmetry and locally break inversion symmetry. This suggests that a finite SHG detected by the 800-nm probe field cannot be directly used to justify the emergence of ferroelectric order. We therefore call for a careful examination of this signal.
Acknowledgments.—F.Y. and X.J.L. did the theoretical and experimental studies, respectively, and contributed equally to this work. F.Y. and L.Q.C. acknowledge support from the US Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145 as part of Computational Materials Sciences Program. F.Y. and L.Q.C. also appreciate the generous support from the Donald W. Hamer Foundation through a Hamer Professorship at Penn State. X.J.L. and D.T. acknowledge support by the National Science Foundation under Grant No. DMR-1554866.